OpenStudy (anonymous):

In a half life word problem, asking to find an interest compounded annually, what do you plug in for n in the formula?

OpenStudy (wolf1728):

Compound Interest Formula Total Earned = principal x (1 + rate)^n Generally the 'n' is the amount of years the money has been invested and compounded.

OpenStudy (anonymous):

The rate is 6.25% compound annually so I'm unsure what to plug in for n

OpenStudy (wolf1728):

Well how many years is the money invested? If it is $5,000 for 3 years then Total Earned = principal x (1 + rate)^n Total = 5,000 * (1.0625) ^ 3 Total = $5,997.31